# Integrable modules over quantum symmetric pair coideal subalgebras

### v2

@@ 1.1. Representation theory
The quantum groups (a.k.a. quantized enveloping algebras) were introduced by Drinfeld [Dr85] and Jimbo [J85] independently, and their representation theory has been studied for many years and been applied to a number of areas of mathematics and mathematical physics such as orthogonal polynomials, combinatorics, knot theory, and integrable systems. One of the central objects in the representation theory of quantum groups is the integrable modules. The notion of integrable modules originates in the representation theory of semisimple Lie algebras. In this context, an integrable module is a module over a semisimple Lie algebra which can be integrated to a representation of a corresponding Lie group. The integrability can be described as the local nilpotency of the Chevalley generators of the Lie algebra. This description can be used to define the notion of integrable modules over quantum groups as it is.
The $\imath$quantum groups (a.k.a. quantum symmetric pair coideal subalgebras) of finite type were formulated by Letzter [Le99] in order to perform noncommutative harmonic analysis on quantum symmetric spaces based on earlier works such as [NS95], [Di96], [N96]. Letzter’s construction was further generalized to infinite type by Kolb [Ko14].
It has been pointed out that the representation theory of $\imath$quantum groups develops very slowly ([Le19], [Wa21]). In fact, we do not still know any general theory to classify the finite-dimensional simple modules over $\imath$quantum groups of finite type, although such $\imath$quantum groups are quantum deformations of the universal enveloping algebra of a finite-dimensional complex reductive Lie algebra. Several type-dependent classification can be found in e.g. [GK91], [IK05], [M06] [IT10], [Wa20], [We20], [Wa21]. Since the $\imath$quantum groups have close resemblance to the quantum groups, it seems to be essential for better understanding of their representation theory to formulate the notion of integrable modules over $\imath$quantum groups. This is the aim of the present paper.
@@ 1.2. Integrable modules
Let us turn to the $\imath$quantum groups. An $\imath$quantum group is a certain subalgebra of a quantum group $\mathbf{U}$, and has a distinguished family $\{E_{j},F_{j},B_{k}\mid j\in I_{\bullet},\ k\in I_{\circ}\}$ of elements, which play similar roles to Chevalley generators of $\mathbf{U}$ in some sense. Here, the $I_{\bullet}$ is a subset of $I$, and $I_{\circ}$ is its complement $I\setminus I_{\bullet}$. An easy idea to define the integrable modules over the $\imath$quantum group is the local nilpotency with respect to the $E_{j}$, $F_{j}$, and $B_{k}$. However, this is not acceptable since the $B_{k}$ may acts semisimply on many modules which should be included in the class of integrable modules. Bao and Wang, in their theory of $\imath$canonical bases [BW18], [BW21], introduced the notion of $\imath$divided powers, which are analogues of the divided powers of quantum groups. Using the $\imath$divided powers, one may be able to resolve the semisimplicity issue of $B_{k}$ above. Instead, testing integrability would be too hard since the $\imath$divided powers are very complicated in general.
@@
The $\mathbf{U}$-modules ${}^{\omega}V(\lambda)\otimes V(\mu)$ appear in Lusztig’s construction of the canonical basis of modified form of $\mathbf{U}$ ([Lu92]). They are replaced by certain $\mathbf{U}^{\imath}$-modules $L^{\imath}(\lambda,\mu)$ in Bao-Wang’s theory of $\imath$canonical bases [BW18], [BW21]. Hence, it is natural to define an integrable $\mathbf{U}^{\imath}$-module to be locally a quotient of $V^{\imath}(\lambda,\mu)$. This definition turns out to work well in the present paper as summarized below.
@@ 1.3. Results
First, we determine a presentation of the $\mathbf{U}^{\imath}$-modules $V^{\imath}(\lambda,\mu)$. This enables one to test a given $\mathbf{U}^{\imath}$-module to be integrable in a systematical way. In many cases, one only needs to check the local nilpotency with respect to $E_{j}$, $F_{j}$, $B_{k}$ just like the quantum group case. In the remaining case, one needs to investigate local semisimplicity of some $B_{k}$. As an application, we show that each integrable $\mathbf{U}$-module is integrable as a $\mathbf{U}^{\imath}$-module.
Next, we concentrate on the $\imath$quantum groups of finite type. We also assume that the parameters of the $\mathbf{U}^{\imath}$ is chosen in a way such that the $\mathbf{U}^{\imath}$ is invariant under a certain anti-algebra involution $\rho$ on $\mathbf{U}$, and that the $\imath$canonical basis of the modified form of $\mathbf{U}^{\imath}$ is stable (or, strongly compatible). It turns out that each simple integrable $\mathbf{U}^{\imath}$-module is finite-dimensional and appears as submodule of a finite-dimensional $\mathbf{U}$-module.
Let us consider the space of matrix coefficients of all simple integrable $\mathbf{U}^{\imath}$-modules. As in the quantum group case ([Ka93, Section 7]), it admits a Peter-Weyl type decomposition. Furthermore, by using the stability of $\imath$canonical bases and the characterization of integrability, we show that this space has a basis which is dual to the $\imath$canonical basis of the modified form of $\mathbf{U}^{\imath}$. This shows that the space of matrix coefficients coincides with Bao-Song’s quantized coordinate ring [BS22], and hence, gives an intrinsic description of the latter.
@@ Acknowledgments
The author thanks Paul Terwilliger for comments on the representation theory of $\imath$quantum groups, especially the $q$-Onsager algebra. He is also grateful to Stefan Kolb for fruitful discussion on a sufficient condition for a weight $\mathbf{U}^{\imath}$-module to be integrable, which stimulates him to formulate Proposition 4.3.1. This work was supported by JSPS KAKENHI Grant Number JP24K16903.
@@ 2. Quantum groups
In this section, we prepare necessary notation and fundamental results regarding the quantum groups. We more or less follow Lusztig’s textbook [Lu93] except in Subsection 2.4, where we discuss Levi and parabolic subalgebras of a quantum group.
@@ 2.1. Cartan and root data
Let $\mathrm{Br}$ and $W$ denote the braid group associated with $I$, and $s_{i}$ the simple reflection corresponding to $i\in I$. They act on $\mathbb{Z}[I]$, $Y$, and $X$ by
@@
Let $\leq$ denote the dominance order on $X$, the partial ordering defined as follows: For each $\lambda,\mu\in X$, we have $\lambda\leq\mu$ if and only if $\mu-\lambda\in Q^{+}$.
@@ 2.2. Quantum groups
Let $\mathbf{U}$ denote the quantum group associated with the root datum $(Y,X,\langle,\rangle,\Pi^{\vee},\Pi)$. Namely, the $\mathbf{U}$ is the unital associative $\mathbb{Q}(q)$-algebra with generators
@@
The braid group $\mathrm{Br}$ acts on $\mathbf{U}$ by $s_{i}\mapsto T^{\prime\prime}_{i,1}$ ([Lu93, §37.1]). For each $w\in W$ with reduced expression $w=s_{i_{1}}\cdots s_{i_{l}}$, set
@@
Let $\mathbf{B}(\pm\infty)$ denote the canonical basis of $\mathbf{U}^{\mp}$. They are graded by $Q^{\mp}$:
@@
For each $b\in\mathbf{B}(\pm\infty)$, we set $\operatorname{wt}(b):=\lambda$.
@@ A $\mathbf{U}$-module $V$ is said to be a weight module if it admits a linear space decomposition
is a $\mathbf{U}$-module isomorphism ([Lu93, 23.3.1 (c)]). The $\mathbf{U}$-module $M(\lambda,\mu)$ (or rather $\dot{\mathbf{U}}1_{-\lambda+\mu}$) is a universal weight module of weight $-\lambda+\mu$ in the following sense: For each weight $\mathbf{U}$-module $V$ and a weight vector $v\in V_{-\lambda+\mu}$, there exists a unique $\mathbf{U}$-module homomorphism $M(\lambda,\mu)\rightarrow V$ which sends $m_{\lambda,\mu}$ to $v$.
@@
$\mathbf{B}(-\infty)[-\lambda]:=\{b\in\mathbf{B}(-\infty)\mid b\cdot{}^{\omega}v_{\lambda}\neq 0\}\rightarrow{}^{\omega}\mathbf{B}(\lambda).$
@@
The $\mathbf{U}$-modules $V(\lambda,\mu)$ are universal integrable $\mathbf{U}$-modules in the following sense: Let $V$ be an integrable $\mathbf{U}$-module and $v\in V$ a weight vector. Then, there exist $\lambda,\mu\in X^{+}$ and a $\mathbf{U}$-module homomorphism $V(\lambda,\mu)\rightarrow V$ which sends $v_{\lambda,\mu}$ to $v$ ([Lu93, Proposition 23.3.10]).
@@ 2.4. Levi and parabolic subalgebras
The $\mathbf{L}$ itself is the quantum group associated with the root datum of type $J$ above. Let $\mathbf{L}^{-}$ and $\mathbf{L}^{+}$ denote the negative and the positive parts of $\mathbf{L}$, respectively. Noting that the Cartan part of $\mathbf{L}$ is the $\mathbf{U}^{0}$, we have a triangular decomposition
@@
Let $\mathbf{P}^{+}=\mathbf{P}^{+}_{J}$ and $\mathbf{P}^{-}=\mathbf{P}^{-}_{J}$ denote the parabolic and the opposite parabolic subalgebra of $\mathbf{U}$ associated with the subset $J$. Namely, the $\mathbf{P}^{+}$ (resp., $\mathbf{P}^{-}$) is the subalgebra of $\mathbf{U}$ generated by $\mathbf{L}$ and
$\{E_{k}\mid k\in K\}\ (\text{resp., $\{F_{k}\mid k\in K\}$}).$
@@
Let $\mathbf{R}^{+}=\mathbf{R}^{+}_{J}$ (resp., $\mathbf{R}^{-}=\mathbf{R}^{-}_{J}$) denote the nilradical part of $\mathbf{P}^{+}$ (resp., $\mathbf{P}^{-}$), that is, the two-sided ideal of $\mathbf{U}^{+}$ (resp., $\mathbf{U}^{-}$) generated by
$\{E_{k}\mid k\in K\}\ (\text{resp., $\{F_{k}\mid k\in K\}$}).$
Then, we have
@@
To prove of the second assertion, let us compute as
@@
The linear isomorphism (2.4.9) has appeared in [BW21, §3.5].
@@
That these maps are well-defined can be straightforwardly verified. Clearly, we have $\phi\circ\psi=\mathrm{id}$ and see that $\psi$ is a $\mathbf{U}$-module homomorphism. Thus, we complete the proof. ∎
@@ 3. Quantum symmetric pairs
After formulating basic notation regarding quantum symmetric pairs following mainly [BW21], we will introduce the notion of integrable modules over an $\imath$quantum group.
@@ 3.1. Quantum symmetric pair
Let $(I,I_{\bullet},\tau)$ be a generalized Satake diagram ([RV20, Definition 1]). From now on, we assume that the root datum $(Y,X,\langle,\rangle,\Pi^{\vee},\Pi)$ is a Satake datum of type $(I,I_{\bullet},\tau)$ ([Wa23c, Definition 3.1.3]). In particular, the $I_{\bullet}$ is a subdatum of $I$ which is supposed to be of finite type, the $\tau$ is an involutive automorphism on the Cartan datum $I$, and the lattices $Y$ and $X$ are supposed to be equipped with involutive automorphisms $\tau$ which are compatible with the one on $I$. Let $w_{\bullet}\in W$ denote the longest element of the Weyl group $W_{\bullet}\subseteq W$ associated with $I_{\bullet}$. Set
@@
For later use, let us recall one of the axioms for generalized Satake diagrams from ([RV20, (2.19)]):
@@
Our assumption on $k$ that $w_{\bullet}k\neq k$ implies this value cannot be 00.
@@
By axiom (3.1.1), the rightmost-hand side cannot be $-1$. Hence, the assertion follows. ∎
@@
They are as general as possible (see [Ko14, Definition 5.6]).
@@ 3.2. Weight modules
A $\mathbf{U}^{\imath}$-module $V$ is said to be a weight module ([Wa23a, Definition 3.3.2]) if it admits a linear space decomposition
@@
as $\mathbf{U}$-modules. The $\mathbf{U}^{\imath}$-module $M^{\imath}(\lambda,\mu)$ is a universal weight module of weight $\overline{w_{\bullet}\lambda+\mu}$ in the following sense:
@@
By isomorphism (3.2.2) and Proposition 3.1.2, we have
@@
as $\mathbf{U}^{\imath}$-modules. Then, the assertion can be found in [Wa23a, after Definition 3.3.5]. ∎

### v3

@@ 1.1. Representation theory
The quantum groups (a.k.a. quantized enveloping algebras) were introduced by Drinfeld [5] and Jimbo [9] independently, and their representation theory has been studied for many years and been applied to a number of areas of mathematics and mathematical physics such as orthogonal polynomials, combinatorics, knot theory, and integrable systems. One of the central objects in the representation theory of quantum groups is the integrable modules. The notion of integrable modules originates in the representation theory of semisimple Lie algebras. In this context, an integrable module is a module over a semisimple Lie algebra which can be integrated to a representation of a corresponding Lie group. The integrability can be described as the local nilpotency of the Chevalley generators of the Lie algebra. This description can be used to define the notion of integrable modules over quantum groups as it is.
The $\imath$quantum groups (a.k.a. quantum symmetric pair coideal subalgebras) of classical type were constructed by using solutions to reflection equation in some papers such as [19], [4], [18] in order to perform noncommutative harmonic analysis on quantum symmetric spaces. Letzter [13] provided a unified construction of $\imath$quantum groups of all finite types (without relying on solutions to reflection equation), and generalized earlier results just mentioned. Her construction was further generalized to infinite types by Kolb [11].
It has been pointed out that the representation theory of $\imath$quantum groups develops very slowly ([14], [23]). In fact, we do not still know any general theory to classify the finite-dimensional simple modules over $\imath$quantum groups of finite type, although such $\imath$quantum groups are quantum deformations of the universal enveloping algebra of a finite-dimensional complex reductive Lie algebra. Several type-dependent classification can be found in e.g. [6], [7], [17], [8], [22], [27], [23], [12]. Since the $\imath$quantum groups have close resemblance to the quantum groups, it seems to be essential for better understanding of their representation theory to formulate the notion of integrable modules over $\imath$quantum groups. This is the aim of the present paper.
@@ 1.2. Integrable modules
Let us turn to the $\imath$quantum groups. An $\imath$quantum group is a certain subalgebra of a quantum group $\mathbf{U}$, and has a distinguished family $\{E_{j},F_{j},B_{k}\mid j\in I_{\bullet},\ k\in I_{\circ}\}$ of elements, which play similar roles to Chevalley generators of $\mathbf{U}$ in some sense. Here, $I_{\bullet}$ is a subset of $I$, and $I_{\circ}$ is its complement $I\setminus I_{\bullet}$. An easy idea to define the integrable modules over the $\imath$quantum group is the local nilpotency with respect to $E_{j}$, $F_{j}$, and $B_{k}$. However, this is not acceptable since $B_{k}$ may act semisimply on many modules which should be included in the class of integrable modules. Bao and Wang, in their theory of $\imath$canonical bases [2], [3], introduced the notion of $\imath$divided powers, which are analogues of the divided powers of quantum groups. Using the $\imath$divided powers, one may be able to resolve the semisimplicity issue of $B_{k}$ above. However, testing integrability would be too hard since the $\imath$divided powers are very complicated in general.
@@
The $\mathbf{U}$-modules ${}^{\omega}V(\lambda)\otimes V(\mu)$ appear in Lusztig’s construction of the canonical basis of modified form of $\mathbf{U}$ ([15]). They are replaced by certain $\mathbf{U}^{\imath}$-modules $L^{\imath}(\lambda,\mu)$ in Bao-Wang’s theory of $\imath$canonical bases [2], [3]. Hence, it is natural to define an integrable $\mathbf{U}^{\imath}$-module to be locally a quotient of $L^{\imath}(\lambda,\mu)$. This definition turns out to work well in the present paper as summarized below.
@@ 1.3. Results
First, we determine a presentation of the $\mathbf{U}^{\imath}$-modules $L^{\imath}(\lambda,\mu)$ (Theorem 4.2.6). This enables one to test if a given $\mathbf{U}^{\imath}$-module is integrable in a systematic way. In many cases, one only needs to check the local nilpotency with respect to $E_{j}$, $F_{j}$, $B_{k}$ just like the quantum group case. In the remaining case, one needs to investigate local semisimplicity of some $B_{k}$. As an application, we show that each integrable $\mathbf{U}$-module is integrable as a $\mathbf{U}^{\imath}$-module (Proposition 4.3.2).
Next, we concentrate on the $\imath$quantum groups of finite type. We also assume that the parameters of the $\mathbf{U}^{\imath}$ is chosen in a way such that the $\mathbf{U}^{\imath}$ is invariant under a certain anti-algebra involution $\rho$ on $\mathbf{U}$, and that the $\imath$canonical basis of the modified form of $\mathbf{U}^{\imath}$ is stable (or, strongly compatible). It turns out that each simple integrable $\mathbf{U}^{\imath}$-module is finite-dimensional and appears as a submodule of a finite-dimensional $\mathbf{U}$-module.
Let us consider the space of matrix coefficients of all simple integrable $\mathbf{U}^{\imath}$-modules. As in the quantum group case ([10, Section 7]), it admits a Peter-Weyl type decomposition. Furthermore, by using the stability of $\imath$canonical bases and the characterization of integrability, we show that this space has a basis which is dual to the $\imath$canonical basis of the modified form of $\mathbf{U}^{\imath}$ (Theorem 5.2.4). This shows that the space of matrix coefficients coincides with Bao-Song’s quantized coordinate ring [1], and hence, gives an intrinsic description of the latter.
@@ Acknowledgments
The author thanks Paul Terwilliger for comments on the representation theory of $\imath$quantum groups, especially the $q$-Onsager algebra. He is also grateful to Stefan Kolb for fruitful discussion on a sufficient condition for a weight $\mathbf{U}^{\imath}$-module to be integrable, which stimulates him to formulate Proposition 4.3.1. He would like to express his gratitude to the anonymous referees for valuable comments and suggestions. This work was supported by JSPS KAKENHI Grant Number JP24K16903.
@@ 2. Quantum groups
In this section, we prepare necessary notation and fundamental results regarding the quantum groups. We more or less follow Lusztig’s textbook [16] except in Subsection 2.4, where we discuss Levi and parabolic subalgebras of a quantum group.
@@ 2.1. Cartan and root data
Let $\mathrm{Br}$ and $W$ denote the braid group and Weyl group associated with $I$, respectively, and $s_{i}$ the simple reflection corresponding to $i\in I$. They act on $\mathbb{Z}[I]$, $Y$, and $X$ by
@@
Let $\leq$ denote the dominance order on $X$, the partial ordering defined by declaring $\lambda\leq\mu$ to mean $\mu-\lambda\in Q^{+}$.
@@ 2.2. Quantum groups
Let $\mathbf{U}$ denote the quantum group associated with the root datum $(Y,X,\langle,\rangle,\Pi^{\vee},\Pi)$. Namely, $\mathbf{U}$ is the unital associative $\mathbb{Q}(q)$-algebra with generators
@@
The braid group $\mathrm{Br}$ acts on $\mathbf{U}$ by $s_{i}\mapsto T^{\prime\prime}_{i,1}$ ([16, §37.1]). For each $w\in W$ with reduced expression $w=s_{i_{1}}\cdots s_{i_{l}}$, set
@@
Let $\mathbf{B}(\pm\infty)$ denote the canonical bases of $\mathbf{U}^{\mp}$, respectively. They are graded by $Q^{\mp}$:
@@
For each $b\in\mathbf{B}(\pm\infty)_{\lambda}$, we set $\operatorname{wt}(b):=\lambda$.
@@ A $\mathbf{U}$-module $V$ is said to be a weight module if it admits a linear space decomposition
is a $\mathbf{U}$-module isomorphism ([16, 23.3.1 (c)]). The $\mathbf{U}$-module $M(\lambda,\mu)$ (or rather $\dot{\mathbf{U}}1_{-\lambda+\mu}$) is a universal weight module of weight $-\lambda+\mu$ in the following sense. For each weight $\mathbf{U}$-module $V$ and a weight vector $v\in V_{-\lambda+\mu}$, there exists a unique $\mathbf{U}$-module homomorphism $M(\lambda,\mu)\rightarrow V$ which sends $m_{\lambda,\mu}$ to $v$.
@@
$\mathbf{B}(-\infty)[-\lambda]:=\{b\in\mathbf{B}(-\infty)\mid b\cdot{}^{\omega}v_{\lambda}\neq 0\}\rightarrow{}^{\omega}\mathbf{B}(\lambda):=\{{}^{\omega}b\mid b\in\mathbf{B}(\lambda)\}.$
@@
The $\mathbf{U}$-modules $V(\lambda,\mu)$ are universal integrable $\mathbf{U}$-modules in the following sense. Let $V$ be an integrable $\mathbf{U}$-module and $v\in V$ a weight vector. Then, there exist $\lambda,\mu\in X^{+}$ and a $\mathbf{U}$-module homomorphism $V(\lambda,\mu)\rightarrow V$ which sends $v_{\lambda,\mu}$ to $v$ ([16, Proposition 23.3.10]).
@@ 2.4. Levi and parabolic subalgebras
The algebra $\mathbf{L}$ itself is the quantum group associated with the root datum of type $J$ above. Let $\mathbf{L}^{-}$ and $\mathbf{L}^{+}$ denote the negative and the positive parts of $\mathbf{L}$, respectively. Noting that the Cartan part of $\mathbf{L}$ is $\mathbf{U}^{0}$, we have a triangular decomposition
@@
Let $\mathbf{P}^{+}=\mathbf{P}^{+}_{J}$ and $\mathbf{P}^{-}=\mathbf{P}^{-}_{J}$ denote the parabolic and the opposite parabolic subalgebra of $\mathbf{U}$ associated with the subset $J$, respectively. Namely, $\mathbf{P}^{+}$ is the subalgebra of $\mathbf{U}$ generated by $\mathbf{L}$ and $\{E_{k}\mid k\in K\}$, and $\mathbf{P}^{-}$ generated by $\mathbf{L}$ and $\{F_{k}\mid k\in K\}$.
@@
Let $\mathbf{R}^{\pm}=\mathbf{R}^{\pm}_{J}$ denote the nilradical part of $\mathbf{P}^{\pm}$. That is, $\mathbf{R}^{+}$ is the two-sided ideal of $\mathbf{U}^{+}$ generated by $\{E_{k}\mid k\in K\}$, and $\mathbf{R}^{-}$ generated by $\{F_{k}\mid k\in K\}$. Then, we have
@@
To prove the second assertion, let us compute as
@@
The linear isomorphism (2.4.9) has appeared in [3, §3.5].
@@
That these maps are well-defined can be straightforwardly verified. Clearly, we have $\phi\circ\psi=\mathrm{id}$, and see that $\psi$ is a $\mathbf{U}$-module homomorphism. Thus, we complete the proof. ∎
@@ 3. Quantum symmetric pairs
After formulating basic notation regarding quantum symmetric pairs following mainly [3], we will introduce the notion of integrable modules over an $\imath$quantum group.
@@ 3.1. Quantum symmetric pair
Let $(I,I_{\bullet},\tau)$ be a generalized Satake diagram ([21, Definition 1]). From now on, we assume that the root datum $(Y,X,\langle,\rangle,\Pi^{\vee},\Pi)$ is a Satake datum of type $(I,I_{\bullet},\tau)$ ([26, Definition 3.1.3]). In particular, $I_{\bullet}$ is a subdatum of $I$ which is supposed to be of finite type, $\tau$ is an involutive automorphism on the Cartan datum $I$, and $Y$ and $X$ are supposed to be equipped with involutive automorphisms $\tau$ which are compatible with the one on $I$. Let $w_{\bullet}\in W$ denote the longest element of the Weyl group $W_{\bullet}\subseteq W$ associated with $I_{\bullet}$. Set
@@
For later use, let us recall one of the axioms for generalized Satake diagrams from ([21, (2.19)]):
@@
Our assumption on $k$ that $w_{\bullet}k\neq k$ implies that this value cannot be 00.
@@
By the axiom (3.1.1), the rightmost-hand side cannot be $-1$. Hence, the assertion follows. ∎
@@
They are as general as [11, Definition 5.6].
@@ 3.2. Weight modules
A $\mathbf{U}^{\imath}$-module $V$ is said to be a weight module ([24, Definition 3.3.2]) if it admits a linear space decomposition
@@
as $\mathbf{U}$-modules. The $\mathbf{U}^{\imath}$-module $M^{\imath}(\lambda,\mu)$ is a universal weight module of weight $\overline{w_{\bullet}\lambda+\mu}$ in the following sense.
@@
By the isomorphism (3.2.2) and Proposition 3.1.2, we have
@@
as $\mathbf{U}^{\imath}$-modules. Then, the assertion can be found in [24, after Definition 3.3.5]. ∎